A worker tearing down a barn, uses a pinch bar 4 ft. long with the shorter end 2in. from the fulcrum. If the worker can excert a force of 150 lbs. at one end of the pinch bar, how great a force can he excert at the other end? (Note the differences in units, 2 in. and 4 ft. Units must be the same when using the lever formula.)
w1 d1 = w2d2
I'm lost after this. Can you help?
Dorothy

Found 2 solutions by checkley75, stanbon:
Answer by checkley75(3666)

48*150=2x
2x=7200
x=7200/2
x=3600 pounds of force

Answer by stanbon(48558)

A worker tearing down a barn, uses a pinch bar 4 ft. long with the shorter end 2in. from the fulcrum. If the worker can excert a force of 150 lbs. at one end of the pinch bar, how great a force can he excert at the other end?
w1 d1 = w2d2
--------------------
w1*2 = 150*48
w1 = 3600 lbs. of force
---------------------
Cheers,
Stan H.